Solve the following initial value problem using the Laplace transform method: y + 8y = -3, y(0) = -5.
Answer #1

Given Equation - Yt8y = -3 ,Y(0)=-5 apply laplace toarsform syls)-ylo) + 8y (5) = -3 syls)-(-5) +8 413)--3 y(s) (5+8)=-3-5 y